Use the shell method to find the volume of the solid generated by revolving the region bounded by the given curves and lines about the y -axis.


y =   3 x 2 , y =  0, x =   4


1
Expert's answer
2021-02-25T00:37:43-0500

Shell method formula:

"\int_a^b2\pi xf(x)dx"

where:


"f(x)=3x^2, a = 0, b = 4\\"

Then:

"\int_0^42\pi x*3x^2dx=6\pi\int_0^4x^3dx=1.5\pi x^4\mid_0^4 =384\pi"

Answer: "384\pi"